OpenStudy (anonymous):

Three times the sum of a number and seven is two less than five times the numbe

OpenStudy (anonymous):

wat is the question?

OpenStudy (anonymous):

that is the question

OpenStudy (anonymous):

so u just want to write out that in numbers?

OpenStudy (aggie12341):

3n+5n-2

OpenStudy (anonymous):

^^wat they said

OpenStudy (aggie12341):

oh @randomaccessmemory .....

OpenStudy (anonymous):

Three times the sum of a number and seven is two less than five times the number. A. 3x + 7 = 5x – 2 B. 3(x + 7) = 5x – 2 C. 3(x + 7) = 2 – 5x D. 3x + 7 = 2 – 5x

OpenStudy (anonymous):

oh wait now that i realize it i think wat aggie said it wrong

OpenStudy (anonymous):

hurry

OpenStudy (anonymous):

so three times the sum of something would mean there are parentheses involved and when u see the word sum that means that you would be adding and when u see less than usually means there will be another part of the equation so the answer would be 3(x+7)=5x-2

OpenStudy (anonymous):

and please do not rush me i am here for your benefit. thank you

OpenStudy (anonymous):

irdgaf!!